OpenStudy (anonymous):

In 2007, the world's population reached 6.7 billion and was increasing at a rate of 1.2% per year. Assume that this growth rate remains constant. (In fact,the growth rate has decreased since 1987.) (a) Write a formula for the world population as a function of the number of years since 2007. (Note: Use the general exponential function.) (b) Use your formula to estimate the population of the world in the year 2032. Round your answer to two decimal places. Ps, I got the answers. My friend did it for me, I just want to know the steps!

OpenStudy (anonymous):

Every year, the population increases by 1.2%. We can write this as a series: \[ a_{n+1} = 0.0012a_n \]We can exant it another term:\[ a_{n+2} = 0.0012a_{n+1} = (0.0012)^2a_n \]The general pattern is: \[ a_{n+r} = (0.0012)^ra_n \]

OpenStudy (anonymous):

Suppose \(n=2007\) and \(r\) is the number or years sense then..

OpenStudy (anonymous):

\[ a_{2007} = 6.7 \]\[ a_{2007+r} = (0.012)^r6.7 \]

OpenStudy (tkhunny):

Aaaa.... NO! Based on that formula, humans will quickly be swept from the planet by the next opportunistic dominant species. Please think about 1.012, rather than 0.012.

OpenStudy (goformit100):

@jaredlol11 "Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."